Pei's savings account balance is $12,000 today. Pei opened the account exactly 7 years ago with a $10,000 deposit. Pei has made no other deposits or withdrawals. What annual interest rate (compounded annually) has the account earned? (Enter your answers as a decimal rounded to 4 decimal places, not a percentage. For example, enter 0.0843 instead of 8.43%)  
Answer #1

Ans 0.0264

FV = Future Value
PV = Present Value
r = rate of interest
n= no of period
FV/ PV = (1 + r )^n
12000 / 10000= (1 + r%)^7
1.20 = (1 + r%)^7
r = 0.0264
Answer #2

SOLUTION :


PV (7 years ago) = 10000 ($)

FV (today) = 12000 ($)

Annual compounding.

Period, n = 7  (annual periods)

Let interest rate be r (in decimals).

So,


FV/PV = (1 + r)^n 

=> 12000/10000 = (1 + r)^7 

=> 1.2 = (1 + r)^7

=> r = (1.2)^(1/7) - 1

=> r = 0.0264 (ANSWER).

answered by: Tulsiram Garg